OpenStudy (anonymous):

solve 4-3v-v=5-3v please

OpenStudy (anonymous):

Combine all the like terms.

OpenStudy (anonymous):

-3v-v = -4v. So 4-4v=5-3v.

OpenStudy (anonymous):

Now add 3v to both sides. 4-v=5.

OpenStudy (anonymous):

Subtract 4 from both sides. -v=1. Divide by -1 to get v positive, and you'll get v=-1.

OpenStudy (anonymous):

You get it?

OpenStudy (anonymous):

Not really..

OpenStudy (anonymous):

You want to combine everything that can go together. You subtract 3v from 4 since 4 has no v attached to it, but you can subtract v from -3v.

OpenStudy (anonymous):

If you do that you'll have -4v. So the equation will read 4-4v=5-3v.

OpenStudy (anonymous):

Now you can either subtract the 4 from the 5 or subtract the 5 from the 4. If you subtract the 4 from the 5 you'll have -4v=1-3v.

OpenStudy (anonymous):

Now add the 3v to the -4v, since there's a - sign in front of the 3v. When you're moving across the equals sign, you always switch the sign infront of the number you're moving. Which is why you subtracted the +4 from the 5 on the other side.

OpenStudy (anonymous):

So now you'll have -v=1. Now divide by -1, since -v is actually -1v. Dividing by -1 will make v positive, which is how teachers almost always want the answer. If you divide both sides by -1 you'll have v=-1.

OpenStudy (anonymous):

Sorry my first comment should say "You CAN'T subtract 3v from 4 since 4 has no v attached......"

OpenStudy (anonymous):

Do you still not get it? I'm not that good at explaining things lol.